Kosumi and Rahquez both leave the internet cafe at the same time, but in opposite directions. If Rahquez travels 5 mph faster than Kosumi and after 8 hours they are 264 miles apart, how fast is each traveling?

Answer by ikleyn(52803)

.

Let x be the slower rate (in miles per hour); then the faster rate is (x+5) mph.


Write an equation for the total separating distance

    8*x + 8*(x+5) = 264.


Simplify step by step

    8x + 8x + 40 = 264

      16x        = 264 - 40

      16x        =   224

        x        =   224/16 = 14.


ANSWER.  The slower rate was 14 mph;  the faster rate was 14 + 5 = 19 mph.


CHECK.  8*14 + 8*19 = 264 miles, the total separating distance.   ! Correct !

Solved.

-----------------

For simple Travel & Distance problems,  see introductory lessons
    - Travel and Distance problems
    - Travel and Distance problems for two bodies moving in opposite directions
    - Travel and Distance problems for two bodies moving in the same direction (catching up)
in this site.

They are written specially for you.

You will find the solutions of many similar problems there.

Read them and learn once and for all from these lessons on how to solve simple Travel and Distance problems.

Become an expert in this area.